Listed below are several fictional stars with their luminosities and distances from Earth given. Rank the stars based on how bright each would appear in the sky as seen from Earth, from brightest to dimmest.  



A) Q-bert: L=400Lsun, d=40ly
B) Frogger: L=400Lsun, d=20ly
C) Pacman: L=200Lsun, d=20ly
D) Centipede: L=200Lsun, d=10ly
E) Galaga: L=100Lsun, d=8ly
Answer #1

just divide given luminosity / square of distance

A) Q-bert: L=400Lsun, d=40ly = 400 / 402 = 0.25
B) Frogger: L=400Lsun, d=20ly = 400 / 202 = 1
C) Pacman: L=200Lsun, d=20ly = 0.5
D) Centipede: L=200Lsun, d=10ly = 2
E) Galaga: L=100Lsun, d=8ly = 1.56

so,

the ranking will be

Centipede > Galaga > Frogger > Pacman > Q-bert